FOURIER TRANSFORM AND INTRACAVITY LASER SPECTROSCOPY OF NICKEL CHLORIDE: IDEN-TIFICATION OF THE X ^{2}Π_{1/2} STATE

\underline{LEAH OBRIEN}, COREY RICE, BARTHEMEAUS OWEN, and TENEIL KELLERMAN, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652; JAMES OBRIEN and HONG CAO, Department of Chemistry, University of Missouri, St Louis, MO 63121-4499.

The near infrared electronic transition of NiCl occurring in the region of 11900 cm^{−1}, also known as System I, has been recorded with rotational resolution by Fourier transform emissions spectroscopy and intracavity laser absorption spectroscopy. This band has been identified as the [12.3] ^{2}Σ^{+} - X ^{2}Π_{1/2} transition. The molecular constants for the newly identified X ^{2}Π_{1/2} state are (in cm^{−1}): T_{0}=385.666(2), B_{0}=0.180778(2), D_{0}=1.652(3)x10^{−7}, p=0.8391(1), p_{D}=3.90(7)x10^{−6}. Results of the analysis will be discussed. The identification of the low-lying X ^{2}Π_{1/2} state completes the analyses of lowlying electronic states of NiCl that correlate to the Ni^{+}3d^{9} electron configuration. Progress on using a 5x5 supermultiplet to describe the low-lying states of NiCl will be reported.